OpenStudy (anonymous):

How many real roots describe the given polynomial function? f(x) = 8x^3 - 2x^2 - 15x

OpenStudy (kagıtucak):

f(x)=x[8x^2-2x-15]=x(4x+5)(2x-3)=0 Thus, x=0, x=-5/4, x=3/2 are the real roots and there are three of them